### v4

# STI-Bench: Are MLLMs Ready for Precise Spatial-Temporal World Understanding?

Abstract
The use of Multimodal Large Language Models (MLLMs) as an end-to-end solution for Embodied AI and Autonomous Driving has become a prevailing trend. While MLLMs have been extensively studied for visual semantic understanding tasks, their ability to perform precise and quantitative spatial-temporal understanding in real-world applications remains largely unexamined, leading to uncertain prospects. To evaluate models’ Spatial-Temporal Intelligence, we introduce STI-Bench, a benchmark designed to evaluate MLLMs’ spatial-temporal understanding through challenging tasks such as estimating and predicting the appearance, pose, displacement, and motion of objects. Our benchmark encompasses a wide range of robot and vehicle operations across outdoor, indoor, and desktop scenarios. The extensive experiments reveal that the state-of-the-art MLLMs still struggle in real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Paper Page: https://mint-sjtu.github.io/STI-Bench.io/
Figure 1: We evaluate state-of-the-art MLLMs on STI-Bench for precise and quantitative spatial-temporal understanding using video inputs. Results indicate the significant challenge in all tasks.
Figure 2: Overview of STI-Bench. We selected the most representative videos from each dataset scene and provided a few simple questions for demonstration.
1 Introduction
The rapid development of Multimodal Large Language Models (MLLMs) [1, 12, 4, 37, 38, 39, 29, 35, 50, 44] has propelled them to the research forefront as a versatile tool to deal with numerous vision and multimodal tasks. Impressive performances have been achieved by MLLMs in general Visual Question Answering tasks [3], which mainly focus on the 2D visual perception and semantic question answering [18, 25, 40, 41, 19, 52, 21].
Beyond 2D visual perception, it has become a prevailing trend to employ MLLMs as an end-to-end solution for Embodied AI [23, 32, 26, 46, 15, 9, 8, 7] and Autonomous Driving [43, 48, 22, 42]. Such tasks require MLLMs to understand the 3D space and time, and then predict optimal manipulation strategies for robotic and vehicular systems. Although many explorations have been conducted, the question remains: Are MLLMs ready for precise spatial-temporal world understanding?
To answer this question, we propose a Spatial-Temporal Intelligence Benchmark (STI-Bench), designed to evaluate MLLMs’ spatial-temporal world understanding capability. We evaluate MLLMs using single video or multiple images as input instead of 3D point clouds. The main reasons are: 1) the majority of state-of-the-art models, e.g., GPT-4o [33] and Gemini [38], can accept images or video as input rather than 3D point clouds; 2) Videos are more frequently used in human’s daily life and they usually contain sufficient information to infer the spatial-temporal environment. STI-Bench contains 300 videos and more than 2,000 QA pairs, covering three major scenarios: Desktop, Indoor, and Outdoor. The videos are sourced from Waymo [36], ScanNet [16], and Omni6DPose [51] respectively, thus encompassing a broad spectrum of real-world environments. As illustrated in Figure 2, we design eight distinct tasks to evaluate models’ ability of static spatial measurement and grounding, and dynamic tasks including speed, acceleration and trajectory estimation.
Through extensive experiments as illustrated in Figure 1, we observe that even the most advanced MLLMs struggle with real-world spatial-temporal understanding, especially in tasks requiring precise distance estimation and motion analysis. Our error analysis reveals three fundamental limitations: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal grounding and integration.
These insights highlight the significant challenges MLLMs face in precisely understanding spatial-temporal information from videos. We believe STI-Bench will serve as an important touchstone that guides the community to distinguish and develop better MLLMs for Embodied AI, Autonomous Driving tasks and beyond.
In summary, our main contributions include:
  • We present STI-Bench, comprising over 300 videos and more than 2,000 tailored questions across outdoor, indoor, and desktop scenarios, providing a systematic quantitative assessment of MLLMs’ spatial-temporal understanding capabilities.
  • We conduct an in-depth study of state-of-the-art video-based MLLMs on STI-Bench, identify key error patterns in spatial-temporal reasoning, and provide empirical insights that can help the community develop more reliable MLLMs for embodied applications.
Benchmark QA Pairs Data Env. Scene View Evaluation Spatial-Temporal D I O Ego Allo. Num. Desc. Dist. Dir. Vel. Traj. SAT [34] 218k I S VSI-Bench [45] 5,156 V R EmbSpatial-Bench [17] 3,640 I R EmbodiedAgentInterface [26] 448 - S - - EmbodiedEval [15] 328 I/V S - - EmbodiedBench [46] 1,128 I S - - WorldSense [6] 3,172 V R MLVU [52] 3,102 V R Video-MMMU [21] 300 V S STI-Bench 2,064 V R
Table 1: Comparison of STI-Bench with existing benchmarks. Data represents the source of our QA data, where V stands for Video and I stands for Image. Env. indicates the environment in which the data is generated, where S represents Simulation and R represents Real. The two columns under View indicate whether the dataset includes Ego-centric and Allocentric perspectives. The two columns under Evaluation specify whether the ground truth is presented in numerical or textual form. The four columns under Spatial-Temporal indicate whether the benchmark evaluates spatial distance, direction (with angular precision), velocity, or a precise and comprehensive trajectory description.
2 Related Work
2.1 Multimodal Large Language Models
MLLMs have achieved groundbreaking performance in visual understanding [1, 4, 12, 38], leveraging large language models (LLMs) [44, 39, 37] and visual encoders. Beyond image-based MLLMs, recent advancements have extended multimodal learning to video understanding. Classical works include models like VideoChat[24], which enable interactive video-based dialogue by integrating multimodal understanding. Subsequent models like Subsequent models like Video-LLaVA[30] enhance visual-language alignment through large-scale vision-language pretraining and fine-tuned adaptation, extending LLaVA[31]’s capability to process video inputs effectively. Recent works, Qwen2.5-VL [44] excels in long-video understanding and temporal localization by incorporating absolute temporal encoding, enabling the model to capture relationships among video frames more effectively. Additionally, its advancement in dynamic resolution modeling allows for seamless adaptation to videos with varying sampling rates, enhancing its versatility in processing diverse video inputs.
2.2 Spatial Understanding with MLLMs
Video MLLMs have attached great importance on semantic understanding. However, spatial understanding has always been a significant challenge, inspiring recent contribution [9, 11, 14]. This progress represents a significant step toward developing world models and embodied agents. Recent advancements in embodied intelligence have explored integrating large-scale MLLMs into robotic control, enabling better generalization and semantic reasoning. RT-2 [8] introduces a vision-language-action framework that transfers web-scale knowledge to robotic control by representing actions as tokens alongside visual and language data, allowing robots to generalize to novel objects and infer multi-step reasoning tasks. Building on this idea, GR-2 [10] extends generalist robot control across diverse embodiments using a Transformer-based architecture trained on a wide range of robotic tasks, demonstrating adaptability across different platforms. Further refining this approach, $\pi_{0}$ [7] incorporates a flow-matching mechanism to generate continuous, precise action trajectories, enhancing fine-grained manipulation skills. By integrating pretrained MLLMs with an independent action module, $\pi_{0}$ achieves zero-shot task execution and flexible adaptation through fine-tuning. Together, these models highlight the potential of leveraging large-scale learning for robotic control, pushing the boundaries of generalization, task adaptability, and multi-modal reasoning in embodied AI.
2.3 Video Benchmarks for MLLM
Recently, multiple benchmarks [41, 40, 52, 19] have emerged for comprehensively evaluating MLLMs’ ability of (long) video understanding, especially about visual perception and semantic reasoning in the form of Video Question Answering. LongVideoBench [41] and LVBench [40] focus on the understanding of long videos. Recent published benchmarks like Video-MME [19] and MMBench-Video [18] comprehensively evaluates MLLMs across various video-related tasks. Existing benchmarks primarily focus on high-level semantic understanding, such as entity recognition and event understanding. In addition, they are largely confined to a temporal extension of 2D image understanding, lacking precise 3D spatial and temporal reasoning of physical quantities. Recent works such as VSI-Bench [45], have shed light on a deeper understanding of the natural world by introducing visual-spatial intelligence tasks for MLLMs, where models are required to provide numerical answers in certain scenarios. However, as illustrated in Table 1, the limited inclusion of scenes and spatial-temporal tasks restricts their ability to capture the complexities of the real physical world. In contrast, STI-Bench comprehensively evaluate models’ ability of precise spatial-temporal understanding in tasks of static spatial measurement and physically motion understanding in Desktop, Indoor and Outdoor scenarios.
Figure 3: Benchmark Statistics. Top: Video length distribution across different categories and datasets. Bottom: The number of questions contributed by each dataset for evaluating different capabilities.
3 STI-Bench
In this section, we present the detailed design and construction of STI-Bench. The construction pipeline is depicted in Figure 4.
Figure 4: Benchmark curation pipeline. The pipeline first aggregates multi-scene RGB datasets that contain 3D bounding box annotations, camera parameters, and point cloud data, which serve as the basis for computing ground truth. From these datasets, we extract numerical ground truth such as distance and velocity, as well as textual descriptions of trajectories and directions. Subsequently, we leverage GPT to assist in generating QA pairs and design a website for rigorous human verification and filtering.
3.1 Task Definition
We propose eight tasks in total, each one systematically examining a distinct aspect of MLLMs’ spatial-temporal understanding. We divide these tasks into two main categories: Static Understanding and Dynamic Understanding.
Static Understanding
  1. a.
    Dimensional Measurement. Concerns estimates of an object’s geometric size, such as length, width, and height, as well as the distance between objects or between the camera and an object. This requires the ability to transform 2D pixel observations into physical world measurements and accurately perceive depth from monocular inputs.
    Example: ”What is the height of this box?” or ”How close is the camera to the table?”
  2. b.
    Spatial Relation. Focuses on identifying spatial relationships among objects or between the camera and an object, including front and back, left and right, up and below. This task tests models’ ability to understand relative positioning across different reference frames and maintain spatial relationship judgment consistency across varying viewpoints.
    Example: ”Is the chair on the left or right side of the table?” or ”What is the position of the red bag relative to the fur sofa?”
  3. c.
    3D Video Grounding. Given a semantic description such as ”the red backpack on the brown sofa,” the goal is to retrieve the object’s 3D bounding box in the camera coordinate system at a specific point in the video. This requires seamlessly aligning linguistic descriptions with visual features and accurately parameterizing 3D positional information.
    Example: ”Locate the 3D bounding box of the red suitcase near the bed.”
Dynamic Understanding
  1. d.
    Displacement and Path Length. Focuses on how far an object or the camera travels between two given time points. This requires tracking consistent reference points across frames and integrating motion information from discrete frames into continuous paths.
    Example: ”How far has the car traveled from 1s to 18s?”
  2. e.
    Speed and Acceleration. Investigates motion parameters by integrating spatial displacement with time intervals. This tests models’ ability to compute spatial derivatives with respect to time and maintain scale consistency across varying distances and perspectives.
    Example: ”What is the average speed of the camera?” or ”How quickly is the ball accelerating?”
  3. f.
    Ego-Centric Orientation. Examines how the camera’s azimuth orientation, parallel to the ground plane, changes over the duration of the video. This requires understanding rotation representations and utilizing fixed scene elements as angular reference points.
    Example: ”How many degrees does the camera’s horizontal orientation shift from the start of the video to its end?”
  4. g.
    Trajectory Description. Describes or infers the camera’s or an object’s motion path throughout the entire video, potentially involving multiple segments of travel and turns. This tests the ability to segment complex trajectories into meaningful components and abstract spatial motion patterns into concise language descriptions.
    Example: ”Summarize the camera trajectory, including distances moved and turns made.”
  5. h.
    Pose Estimation. Given the camera’s initial 3D pose, including position and orientation, estimates its pose at a specified point in the video using only the observed RGB data. This requires visual odometry capabilities and the ability to manage cumulative error in long sequences.
    Example: ”Given the initial pose of the camera, what is the camera’s pose at the requested time?”
Each of these tasks presents unique challenges that collectively evaluate models’ comprehensive spatial-temporal intelligence across different scales, from millimeter-precision desktop manipulation to meter-scale indoor navigation and beyond. Success in these tasks requires not only fundamental 3D spatial reasoning but also physical common sense and the ability to integrate information across different modalities and reference frames over time.
3.2 Benchmark Construction
Data Collection.
To encompass a broad spectrum of real-world environments, STI-Bench covers three major scenarios: Desktop, Indoor, and Outdoor. Accordingly, we draw from three publicly available datasets—Waymo [36] for autonomous driving, ScanNet[16] for indoor 3D scene reconstruction, and Omni6DPose[51] for desktop-scale 6D object pose estimation. These datasets provide frame-by-frame camera intrinsic and extrinsic parameters, as well as point clouds for each object, which we map to two-dimensional bounding boxes in each frame.
Automatic QA Pair Generation.
We used MLLMs to produce detailed semantic descriptions for each object, such as “A beige minivan with a roof rack,” “A refrigerator with emoji magnets, photos, and a to-do list,” or “A red backpack on a brown leather sofa.” Next, leveraging the frame-by-frame annotations, we computed the ground-truth information required for each task. We then provided the ground-truth data, object descriptions, and task-specific QA requirements to MLLMs to generate a diverse set of questions and challenging answer options.
Human Quality Control.
During QA pair generation, several issues arose:
  1. 1.
    LLM-generated descriptions could be inaccurate or fail to uniquely identify the target object.
  2. 2.
    Some questions and options remained unreasonable or incorrect, even with detailed guidelines.
  3. 3.
    In certain cases, the video alone did not provide sufficient information. For example, the camera was occluded but lidar data were available.
To address these challenges, we developed a website for multiple rounds of manual filtering and sampling-based review, ensuring high-quality questions. We also randomly shuffled the answer options to enhance evaluation robustness. Ultimately, we curated more than 2,000 high-quality QA pairs from over 300 videos. Details are shown in Figure 3.
Fine-Grained Adjustment.
To ensure task diversity and challenge, each question is equipped with carefully constructed distractor options that are both significantly different from the correct answer while remaining within a reasonable range. Considering the varying precision requirements across different scenarios—millimeter-level for desktop applications, centimeter-to-decimeter-level for indoor environments, and decimeter-to-meter-level for outdoor scenes—we applied scenario-specific scaling factors to generate distractors and employed logarithmic sampling techniques to distribute numerical differences more evenly, thereby enhancing the robustness of our evaluation.
Specifically, given a scene type $S\in\{\text{Desktop},\text{Indoor},\text{Outdoor}\}$, a correct answer $A_{correct}$, and four initial distractor options $\{A_{d1},A_{d2},A_{d3},A_{d4}\}$, our fine-grained adjustment process consists of the following steps:
[…]
2. We logarithmically sample an error value $e$ from the corresponding error range:
$e=E_{min}\cdot(E_{max}/E_{min})^{u}$
where $u\sim\mathcal{U}(0,1)$ is a uniform random number.
3. We adjust each distractor option $A_{di}$ to $A_{di}^{\prime}$ using a weighted average:
$A_{di}^{\prime}=(1-w)\cdot A_{di}+w\cdot A_{correct}$
where the weight $w$ is identical for all distractors and is determined by solving:
$\min_{i\in\{1,2,3,4\}}\|A_{di}^{\prime}-A_{correct}\|=e$
This yields:
$w=1-\frac{e}{\min_{i\in\{1,2,3,4\}}\|A_{di}-A_{correct}\|}$
This approach ensures that the minimum distance between any adjusted distractor and the correct answer equals the sampled error value $e$, while other distractors maintain relatively larger distances but are adjusted using the same weight factor.
  Methods Rank Avg. Static Understanding Dynamic Understanding
Dim. Meas. Spatial Relation 3D Video Grounding Disp. & P.L. Speed & Acc. Ego Orient. Traj. Desc. Pose Est.
Proprietary Models (API)
GPT-4o[33] 8 34.8 27.1 51.8 29.0 23.2 35.4 33.7 32.0 53.6
Gemini-2.0-Flash[38] 4 38.7 31.9 50.0 31.8 27.7 32.1 10.8 38.5 61.3
Claude-3.7-Sonnet[2] 3 40.5 29.8 45.5 35.7 28.9 38.8 40.0 47.4 62.6
Gemini-2.5-Pro[20] 1 41.4 38.7 53.8 36.9 33.9 33.1 52.5 47.4 50.4
Open-source Models
MiniCPM-V-2.6[47] 10 26.9 27.7 44.5 29.0 19.0 25.7 7.0 30.8 35.6
VideoChat-R1[28] 9 32.8 23.2 47.3 31.5 22.4 31.1 26.0 47.9 48.3
VideoLLaMA3-7B[49] 7 35.2 29.4 48.6 36.1 21.5 36.7 23.2 54.6 48.1
VideoChat-Flash[27] 6 36.3 33.6 51.4 33.1 27.1 32.3 22.2 54.2 51.4
InternVL2.5-78B[13] 5 38.5 29.9 52.8 31.6 24.9 37.2 49.2 43.6 53.6
Qwen2.5-VL-72B[5] 2 40.7 31.5 47.6 39.1 25.1 38.4 43.8 51.3 60.6
Table 2: Evaluation on STI-Bench. Orange = best; Light Orange = second best.Note: The random baseline for all tasks is 20%.
Model Outdoor Indoor Desktop Overall
MiniCPM-V-2.6[47] 27.9 26.1 26.6 26.9
VideoChat-R1_7B[28] 41.1 28.0 27.1 32.8
GPT-4o[33] 41.4 33.1 27.3 35.2
VideoLLaMA3-7B[49] 41.2 31.9 30.5 35.2
Gemini-2.0-Flash[38] 39.1 35.9 30.5 36.0
VideoChat-Flash[27] 38.6 36.7 31.3 36.3
InternVL2.5-78B[13] 46.1 34.5 32.1 38.5
Claude-3.7-Sonnet[2] 47.2 38.2 32.3 40.5
Qwen2.5-VL-72B[5] 50.6 35.0 33.5 40.7
Gemini-2.5-Pro[20] 48.7 37.1 35.8 41.4
Table 3: Evaluation results across different scenes.
Figure 5: Evaluation results across different scenes and tasks.
4 Experiments
4.1 Settings
We conduct a thorough evaluation of leading MLLMs from diverse model families, focusing on both proprietary and open-source solutions. Specifically, we assess the performance of four proprietary models, GPT-4o[33], Gemini-2.0-Flash[38], Gemini-2.5-Pro[38], and Claude-3.7-Sonnet[2], as well as several representative open-source MLLMs that have undergone specialized video-related training, including Qwen2.5-VL-72B[5], InternVL2.5-78B[13], MiniCPM-V-2.6[47], VideoChat-Flash[27],VideoChat-R1[28] and VideoLLaMA3-7B[49].
Considering the stability of open-source models and the API limitations of proprietary models, we uniformly sample 30 frames from the video for each record and explicitly indicate the sampling FPS for the current video within the prompt. An exception is made for Claude-3.7-Sonnet, for which only 20 frames are sampled due to its API constraints. Our benchmark tasks are presented in a multiple-choice format with five possible answers, hence a random guess baseline yields a 20% accuracy. We measure each model’s accuracy by directly comparing the model’s selected answer with the ground truth, without employing any additional external models or annotations for performance evaluation.
Figure 6: The simplified thought process examples of Gemini 2.5 Pro.
4.2 Main Results
As shown in Table 2 and Table 3, we present a comprehensive evaluation of various MLLMs on STI-Bench. Overall, Gemini-2.5-Pro achieves the highest average accuracy of 41.4% among all tested models, closely followed by Qwen2.5-VL-72B at 40.7% among open-source models. While these results significantly exceed the random guess baseline (20%), they still highlight substantial room for improvement in spatial-temporal understanding.
When analyzing performance across different scene types (Table 3), we observe varied patterns. Qwen2.5-VL-72B demonstrates notably strong performance in outdoor scenarios (50.6%), while Gemini-2.5-Pro also performs well outdoors (48.7%). In indoor environments, Claude-3.7-Sonnet (38.2%) and Gemini-2.5-Pro (37.1%) lead. For desktop scenes, Gemini-2.5-Pro shows the strongest performance (35.8%), followed by Qwen2.5-VL-72B (33.5%). These variations suggest different model specializations, possibly influenced by training data distribution or architectural differences in handling diverse environmental cues.
Task-specific performance (Table 2) reveals particularly challenging areas. Models generally struggle with tasks requiring precise quantitative estimation. For instance, the best performance on Dimensional Measurement is achieved by Gemini-2.5-Pro (38.7%), and on Displacement & Path Length also by Gemini-2.5-Pro (33.9%). Speed & Acceleration also presents difficulties, with Claude-3.7-Sonnet leading at 38.8%. In contrast, models demonstrate stronger capabilities in Pose Estimation (best: 62.7% by Claude-3.7-Sonnet) and Spatial Relation tasks (best: 53.8% by Gemini-2.5-Pro). Notably, Gemini-2.5-Pro also achieves the highest score in Ego-Centric Orientation (52.5%) among proprietary models, while VideoLLaMA3-7B leads in Trajectory Description (54.6%) among open-source models.
Among open-source models, Qwen2.5-VL-72B stands out with an average accuracy of 40.7%, demonstrating highly competitive performance. However, other open-source models like InternVL2.5-78B (38.5%) also show promise, though smaller models like MiniCPM-V-2.6 (26.9%) still lag behind their larger counterparts and proprietary models.
It is important to emphasize that even the best-performing model, Gemini-2.5-Pro, achieves only 41.4% average accuracy on our benchmark. While this is considerably better than random guessing, it remains far from the reliability required for demanding real-world applications like embodied AI or autonomous driving. These results indicate that current MLLMs, despite their impressive general visual understanding capabilities, still require significant advancements in precise spatial-temporal intelligence for embodied tasks.
4.3 Experimental Analysis
Given that Gemini-2.5-Pro exhibits strong multi-modal reasoning capabilities, provides detailed thinking processes (as exemplified in Figure 6), and achieves the top overall performance (41.4% average accuracy) among all evaluated models, we select it as a representative for in-depth analysis.
Our analysis of Gemini-2.5-Pro reveals several key characteristics of current state-of-the-art MLLMs’ spatial-temporal understanding capabilities. Overall, the model achieves 41.4% average accuracy. Performance varies across scene types: it performs strongest in outdoor scenarios (48.7%), followed by indoor (37.1%) and desktop environments (35.8%). This performance disparity might suggest that the model’s training data or architecture is better attuned to outdoor and larger-scale understanding, or that such environments typically provide clearer visual cues for spatial reasoning.
When examining task-specific performance for Gemini-2.5-Pro (Table 2), we observe its strongest capabilities in tasks involving relative understanding and object state. It achieves high scores in Spatial Relation (53.8%, the best among all models) and Ego-Centric Orientation (52.5%, the highest among proprietary models), followed by Pose Estimation (50.4%). Interestingly, while tasks requiring precise quantitative estimation generally show lower absolute scores across models, Gemini-2.5-Pro performs very competitively and often leads in these areas: it achieves the top scores in Dimensional Measurement (38.7%) and Displacement & Path Length (33.9%). Its score for Speed & Acceleration is 33.1%. This indicates that while the absolute accuracy in precise estimation tasks remains a challenge, Gemini-2.5-Pro demonstrates comparatively strong capabilities in these demanding domains relative to other models.
By leveraging the model’s reasoning process and uniformly sampling approximately 200 error records across each task type and scenario, we categorize its errors into three representative patterns. Figure 7 shows the distribution of error categories.
Figure 7: Distribution of error categories in Gemini-2.5-Pro across our sampled error cases.
Inaccurate Spatial Quantification
The model faces significant challenges in accurately estimating static spatial properties and relationships from visual inputs. These difficulties manifest when estimating object dimensions (size), distances between objects or between camera and objects, and absolute 3D coordinates/bounding boxes at specific time points. These errors stem from a lack of clear visual size references, difficulty distinguishing between numerically close options, and the inherent challenges of inferring metric scale from 2D pixels and estimating depth with monocular cameras. Such limitations directly impact performance in dimensional measurement, spatial relation, and 3D video grounding tasks.
Flawed Temporal Dynamics Understanding
The model struggles to perceive, track, and interpret cross-frame information that changes over time, such as motion and its dynamics. This results in erroneous calculations or descriptions of displacement, path length, speed, acceleration, directional changes (ego-centric or object pose), and overall trajectory shapes. The model particularly struggles with relative motion (distinguishing object motion from camera motion), a problem exacerbated by sparse temporal sampling. These difficulties arise from challenges in integrating information across frames, lack of internal models for physics/kinematics, inability to separate ego-motion from object motion, and information loss due to sparse sampling. These issues manifest in tasks involving displacement and path length, speed and acceleration, ego-centric orientation, trajectory description, and pose estimation (as it changes over time).
Weak Cross-Modal Grounding and Integration
The model fails to properly connect textual queries/instructions with relevant spatial-temporal visual content, or to integrate provided non-visual data (such as initial poses) with visual information. This includes misinterpreting temporal constraints (like ”from 1s to 18s,” ”at the end,” ”the moment of last co-occurrence”), failing to correctly utilize provided initial conditions (e.g., initial camera pose in pose estimation tasks), and incorrectly associating structured data (coordinates, timestamps) with visual elements. These errors stem from deficiencies in parsing structured/natural language instructions and difficulty integrating information from different modalities (text prompts, initial state data, video frames) into a unified reasoning process. This affects all tasks that rely on specific instructions or initial data.
These error patterns highlight that, despite Gemini-2.5-Pro’s strong performance relative to other models, it still faces significant challenges in precise spatial-temporal understanding. Its limitations in quantitative estimation and complex spatial-temporal reasoning indicate that current MLLMs remain far from achieving the reliability required for embodied AI or autonomous driving applications.
5 Conclusion
We introduced STI-Bench, a comprehensive benchmark to assess MLLMs’ spatial-temporal understanding through over 300 real-world videos and 2,000 QA pairs of robot outdoor, indoor, and desktop scenarios, which reveals significant limitations in current MLLMs’ spatial-temporal understanding capabilities, with top-performing models like Gemini-2.5-Pro achieving around 41.4% average accuracy. Models particularly struggle with precise quantitative tasks like dimensional measurement. Our analysis identifies three key weaknesses: inaccurate spatial quantification, flawed temporal dynamics understanding, and weak cross-modal integration. These findings emphasize the substantial gap between current capabilities and the reliability needed for embodied AI and autonomous driving applications. STI-Bench provides a valuable framework for evaluating and improving MLLMs’ ability to understand the physical world—essential for developing the next generation of embodied intelligent systems.
Acknowledgements
We thank the data annotation team at MolarData for their support, as well as the volunteer contributors Yuxin Liu, Junjie Ruan, Lingxi Xia, Zixuan Huang, Tianrui Wan, Qingcheng Wei, Yujie Yao, Shangyang Dong, Zhuofan Zeng, and Yiming Li for their valuable work as human annotators and evaluators.
Appendix A Details for Benchmark Construction
A.1 Dataset Interface
When searching for datasets, we found that apart from RGB frame data, other data formats are inconsistent. For example, we noticed that it does not provide 3D bounding box information, so we could only use it to compute camera displacement, velocity, and other related physical quantities. In contrast, Waymo includes such information, which is why we aim to align different datasets into a unified format. Therefore, we unify different types of data by converting them into frame-by-frame instance-level 3D point clouds or bounding box annotations, along with per-frame camera parameters, enabling us to obtain all the required information and process it in a consistent manner.
A.2 Ground Truth Computation
In STI-Bench, the ground truth annotations are derived from multi-modal data, including frame-by-frame camera intrinsics/extrinsics and precise 3D point cloud annotations. We utilize three publicly available datasets (Waymo, ScanNet, Omni6DPose) to cover different scenarios (outdoor, indoor, desktop). Below, we provide eight benchmark tasks with unified notations and formulas.
A.2.1 Static Understanding
Dimensional Measurement
Let $l_{x}$, $l_{y}$, $l_{z}$ denote the dimensions (length, width, height) of an object along the $x$, $y$, and $z$ axes:
$l_{x}$ $=x_{\max}-x_{\min},$ (1)
$l_{y}$ $=y_{\max}-y_{\min},$
$l_{z}$ $=z_{\max}-z_{\min}.$
Here, $l_{x},l_{y},l_{z}$ represent the object size along each coordinate axis.
If we need the distance between two objects (or between the camera and an object), let $d_{12}$ be the Euclidean distance between their center points:
$d_{12}=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}.$ (2)
Here, $(x_{1},y_{1},z_{1})$ and $(x_{2},y_{2},z_{2})$ are the center coordinates of the two objects.
Spatial Relation
When the difference along one coordinate axis is significantly larger than along others, the sign of that difference determines the spatial relation:
$r_{xy}$ $=\mathrm{sign}(x_{A}-x_{B}),$ (3)
$r_{yz}$ $=\mathrm{sign}(y_{A}-y_{B}),$
$r_{zx}$ $=\mathrm{sign}(z_{A}-z_{B}).$
Here, $r_{xy},r_{yz},r_{zx}$ indicate relative positioning along each axis (e.g., front/back, left/right, above/below). We choose the axis with the greatest difference to label the dominant relation.
3D Video Grounding
For frame $t$ in the camera coordinate system, the 3D bounding box of an object can be described with dimensions, center position, and optional rotations:
$\text{BBox}_{t}=(\,l_{t},\,w_{t},\,h_{t},\,x_{t},\,y_{t},\,z_{t},\,\text{yaw}_{t},\,\text{pitch}_{t},\,\text{roll}_{t}\,).$ (4)
Here, $(l_{t},w_{t},h_{t})$ are the object dimensions, $(x_{t},y_{t},z_{t})$ is the center position, and $(\text{yaw}_{t},\text{pitch}_{t},\text{roll}_{t})$ are optional rotation angles if available.
A.2.2 Dynamic Understanding
Pose Estimation
Given the camera’s initial pose $(p_{0},o_{0})$, the pose $(p_{t},o_{t})$ at time $t$ can be obtained using the extrinsic-derived matrices $R_{t}$ (rotation) and $T_{t}$ (translation):
$p_{t}$ $=R_{t}\,p_{0}+T_{t},$ (5)
$o_{t}$ $=o_{0}+\Delta o_{t}.$
Here, $p_{t}$ is the position, $o_{t}$ is the orientation.
Displacement and Path Length
Let $p_{i}=(x_{i},y_{i},z_{i})$ be the position at time $i$. The displacement $d_{0n}$ and path length $L_{\text{traj}}$ are computed as:
$d_{0n}=\sqrt{(x_{n}-x_{0})^{2}+(y_{n}-y_{0})^{2}+(z_{n}-z_{0})^{2}},$ (6)
$L_{\text{traj}}=\sum_{i=1}^{n}\sqrt{(x_{i}-x_{i-1})^{2}+(y_{i}-y_{i-1})^{2}+(z_{i}-z_{i-1})^{2}}.$ (7)
Here, $d_{0n}$ is the straight-line distance from the initial to the final position; $L_{\text{traj}}$ sums consecutive segment lengths for the entire path.
Speed and Acceleration
Let $\Delta t$ be the time interval between consecutive frames. Then the speed $v_{i}$ and acceleration $a_{i}$ are:
$v_{i}$ $=\frac{d_{i}}{\Delta t},$ (8)
$a_{i}$ $=\frac{v_{i}-v_{i-1}}{\Delta t}.$
Here, $d_{i}$ is the displacement between adjacent frames, $v_{i}$ is the speed at time $i$, and $a_{i}$ is the acceleration.
Ego-Centric Orientation
If $\theta_{t}$ denotes the camera orientation (azimuth) at time $t$, then the orientation change $\Delta\theta_{t}$ is:
$\Delta\theta_{t}=\theta_{t}-\theta_{0}.$ (9)
This indicates how much the camera has rotated relative to its initial azimuth.
Trajectory Description
We apply the Ramer-Douglas-Peucker (RDP) algorithm to simplify the sequence of positions into key line segments. The resulting polyline is described in a piecewise manner (e.g., “go straight for 30m, turn left 85°, then go straight for 20m, …”), providing a concise representation of complex motion trajectories.
A.3 Human Involved Quality Control
Despite leveraging automated methods for generating large quantities of question-answer pairs, several critical issues necessitated human intervention for quality assurance. These included inaccuracies or ambiguities in automatically generated object descriptions, logical inconsistencies or irrelevant content within questions and options, and insufficient visual information in video data due to occlusions or limited resolution. To address these challenges, we established a dedicated online platform for systematic multi-round manual review and random sampling checks. Human reviewers evaluated each QA pair based on several criteria: clarity and accuracy of object descriptions, logical coherence between questions and answer options, sufficiency of visual information to accurately respond to questions, and effectiveness of distractors in posing genuine challenges. After passing the rigorous review process, answer options were randomized to prevent bias arising from fixed option ordering. Ultimately, approximately 20% of all generated QA pairs were retained. This meticulous procedure enabled us to select over 2,000 high-quality QA pairs from more than 300 video sequences, significantly enhancing the robustness and reliability of the STI-Bench benchmark dataset.
Appendix B Adjustment Details
To ensure option diversity under different positioning accuracies, we generate refined distractors for each scene $S\in\{\mathrm{desktop},\mathrm{indoor},\mathrm{outdoor}\}$. Given the ground-truth answer $A_{\text{c}}$ and four initial distractors $\{A_{d,1},A_{d,2},A_{d,3},A_{d,4}\}$, the adjustment consists of three steps.
Step 1: Scene-dependent error interval.
Define $[E_{\min},E_{\max}]$ as
[…]
Step 2: Log-uniform sampling of target error.
Sample $u\sim\mathcal{U}(0,1)$ and compute
$e=E_{\min}\Bigl{(}\tfrac{E_{\max}}{E_{\min}}\Bigr{)}^{u}.$ (10)
Step 3: Joint adjustment of distractors.
All distractors are shifted towards $A_{\text{c}}$ with a common weight $w$:
$A_{d,i}^{\prime}=(1-w)A_{d,i}+wA_{\text{c}},\qquad i=1,\dots,4.$ (11)
The weight $w$ is chosen such that the closest adjusted distractor attains the target error $e$:
$w=1-\frac{e}{\displaystyle\min_{i}\bigl{\lVert}A_{d,i}-A_{\text{c}}\bigr{\rVert}}.$ (12)
By sharing the same $w$, the minimum gap between any distractor and the correct answer is exactly $e$, while others remain further apart, thereby preserving option dispersion.
Appendix C Detailed Results
To demonstrate the effectiveness of STI-Bench, we present visualization examples of QA pairs from representative outdoor (see Figure 8), indoor (see Figure 9), and desktop (see Figure 10) scenarios.
To provide a representative analysis of the capability bottlenecks of current SOTA models, we sampled over 200 erroneous reasoning processes from Gemini 2.5 Pro and presented one typical correct and one typical incorrect example (see Figure 11 and Figure 12).
In addition, to more concretely quantify the performance of different models on fine-grained categories, Table 4, Table 5, and Table 6 provide a detailed breakdown of model performance across various fine-grained categories within outdoor, indoor, and desktop scenarios, respectively.
Figure 8: STI-Bench Examples (Outdoor)
Figure 9: STI-Bench Examples (Indoor)
Figure 10: STI-Bench Examples (Desktop)
Figure 11: Thinking example (Correct)
Figure 12: Thinking example (Incorrect)
Table 4: Performance on Fine-grained Outdoor Categories.
  Category (Outdoor) Models Avg.
Gemini- 2.5-Pro Qwen2.5-VL -72B Claude-3.7 -Sonnet InternVL 2.5-78B
  Night 37.2% 43.6% 38.5% 35.9% 38.8%
Day 50.3% 47.9% 44.7% 43.3% 46.6%
Dawn/Dusk 45.3% 50.4% 47.0% 47.4% 47.5%
Rainy 51.1% 55.5% 54.5% 48.6% 52.4%
Clear 51.0% 59.4% 50.0% 55.7% 54.0%
Foggy 71.4% 28.6% 71.4% 71.4% 60.7%
Table 5: Performance on Fine-grained Indoor Categories.
  Category (Indoor) Models Avg.
Gemini- 2.5-Pro Qwen2.5-VL -72B Claude-3.7 -Sonnet InternVL 2.5-78B
  laundry room 25.0% 25.0% 50.0% 50.0% 37.5%
copy/mail room 27.6% 30.0% 30.0% 40.0% 31.9%
bookstore/library 32.0% 24.0% 32.0% 28.0% 29.0%
kitchen 26.7% 29.0% 32.3% 32.3% 30.1%
bedroom/hotel 36.6% 33.9% 42.3% 30.4% 35.8%
office 36.2% 33.3% 37.4% 30.2% 34.3%
hallway 38.5% 42.0% 38.5% 42.0% 40.3%
living room/lounge 39.0% 34.6% 40.3% 29.9% 36.0%
conference room 26.8% 33.9% 38.6% 32.2% 32.9%
apartment 39.3% 48.3% 46.4% 44.8% 44.7%
misc 36.4% 45.5% 54.5% 45.5% 45.5%
bathroom 49.1% 34.2% 34.2% 41.2% 39.7%
storage/basement/garage 30.8% 57.1% 35.7% 57.1% 45.2%
Table 6: Performance on Fine-grained Desktop Categories.
  Category (Desktop) Models Avg.
Gemini- 2.5-Pro Qwen2.5-VL -72B Claude-3.7 -Sonnet InternVL 2.5-78B
  Stone 27.3% 16.7% 33.3% 9.1% 21.6%
Paper 40.0% 37.5% 18.8% 25.0% 30.3%
Leather 23.1% 33.3% 40.0% 26.7% 30.8%
Plastic 32.8% 32.6% 28.6% 29.8% 31.0%
Fabric 40.7% 35.5% 30.0% 35.5% 35.4%
Wood 37.8% 33.7% 34.9% 33.7% 35.0%
Tile 39.6% 38.0% 38.0% 40.0% 38.9%
Concrete 42.9% 33.3% 53.3% 46.7% 44.1%